Solve the following equations.
a) sqrt x - 1 = 3
b) sqrt (x^3) = 8
c) ^3 sqrt (x^2) = 4

Answer by uma(370)

(a) sqrt(x) - 1 = 3
adding 1 to both the sides of the equation,
sqrt(x) - 1 + 1 = 3 = 1
==> sqrt(x) = 4
squaring both the sides,
x = 4^2
==> x = 16.



(b) sqrt(x^3) = 8
Squaring both the sides,
x^3 = 8^2
==> x^3 = 64
==> x* x* x = 4*4*4
==> x = 4


(c) cubic root of x^2 = 4
when we cube both the sides,
x^2 = 4^3
==> x^2 = 64
Taking square root on both the sides we get,
x = +8, -8
X = 8 or x = -8

Good Luck!!!